I also need some help on this question if anyone is able to help.
A clothing manufacturer has some pure silk thread and some thread that is 85% silk. How many kilograms of each must be woven together to make 75kg of a blend that is 96% silk? Thanks for your time.

Answer by ankor@dixie-net.com(22740)

A clothing manufacturer has some pure silk thread and some thread that is 85% silk. How many kilograms of each must be woven together to make 75kg of a blend that is 96% silk?
:
Let x = amt in kg of pure silk thread,
:
The resulting total is given as 75 kg, therefore:
(75-x) = amt in kg of the 85% thread
:
The mixture equation:
1.0x + .85(75-x) = .96(75)
:
1x + 63.75 - .85x = 72; multiplied what's inside the brackets
:
1x - .85x = 72 - 63.75
:
.15x = 8.25
:
x = 8.25/.15
:
x = 55 kg of pure silk thread
Then
75 - 55 = 20 kg of 85% thread
:
:
Check solution
1(55) + .85(20) = .96(75)
55 + 17 = 72